### v2

# Unveiling the jet angular broadening with photon-tagged jets in high-energy nuclear collisions

Abstract
The medium modification of jet substructure in hot and dense nuclear matter has garnered significant interest from the heavy-ion physics community in recent years. Measurements of inclusive jets show an angular narrowing in nucleus-nucleus collisions, while recent CMS results for photon-tagged jets ($\gamma$+jets) suggest evidence of broadening. In this study, we conduct a theoretical analysis of the angular structure of inclusive jets and $\gamma$+jets using a transport approach that accounts for jet energy loss and the medium response in the quark-gluon plasma. We examine the girth modification of $\gamma$+jets in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, achieving satisfactory agreement with recent CMS measurements. We explore the relationship between selection bias and jet kinematics by varying the threshold for $x_{j\gamma}=p_{T}^{\rm jet}/p_{T}^{\gamma}$. Notably, we quantitatively demonstrate that $\gamma$+jets significantly reduce selection bias and can effectively select jets that have been sufficiently quenched in PbPb collisions, which is crucial for capture the jet angular broadening. Additionally, we estimate the contributions of medium-induced gluon radiation and the medium response to the broadening of the jet angular substructure. Lastly, we analyze the modification patterns of jet $R_{g}$ and $\Delta R_{\rm axis}$ in PbPb collisions, which indicate slight broadening for $\gamma$+jets and noticeable narrowing for inclusive jets compared to pp collisions.
pacs:
25.75.Ld, 25.75.Gz, 24.10.Nz
I Introduction
High-energy collisions of heavy nuclei at the Relativistic Heavy Ion Collider (RHIC) and the Large Hadron Collider (LHC) provide an experimental avenue to unravel the mysteries of the quark-gluon plasma (QGP), a short-lived state of de-confined nuclear matter created at extremely high temperature and density. The jet quenching phenomenon, energy dissipation of an energetic parton when passing through the hot and dense nuclear matter, is one of the most important signatures of the QGP formation Gyulassy:2003mc ; Gyulassy:1990ye ; Qin:2015srf ; Vitev:2008rz ; Proceedings:2007ctk ; Vitev:2009rd ; Casalderrey-Solana:2014bpa ; Gyulassy:1993hr ; Wang:2001ifa ; Vitev:2008vk . Investigations on jet quenching reveal the phase structure of the strongly-coupled nuclear matter and push our knowledge of the quantum chromodynamics (QCD) under extreme conditions Connors:2017ptx ; Andrews:2018jcm ; Cao:2020wlm ; Cunqueiro:2021wls ; Apolinario:2022vzg ; Sorensen:2023zkk ; Zhao:2020wcd ; Xie:2024xbn ; Zhang:2021xib ; Tang:2020ame ; Shen:2020mgh ; Chen:2024eaq ; Chen:2024aom ; Zhao:2022dac ; Shou:2024uga .
Jet substructures are valuable tools to gain insight into the details of the jet-medium interaction in the QGP, such as the medium-induced gluon radiation Baier:1996kr ; Baier:1996sk , medium response  KunnawalkamElayavalli:2017hxo ; Pablos:2019ngg ; Chen:2020tbl ; Casalderrey-Solana:2020rsj ; He:2018xjv ; Ke:2020clc , medium resolution length Hulcher:2017cpt ; Mehtar-Tani:2016aco ; Caucal:2018dla and the “Molière elastic scattering” DEramo:2012uzl ; DEramo:2018eoy . Recent reviews on this topic can be found in references Cunqueiro:2021wls ; CMS:2024krd ; Arslandok:2023utm ; Apolinario:2022vzg ; Marzani:2019hun . A significant focus of recent investigations has been on how the angular structure of jets is modified in nucleus-nucleus collisions—whether they narrow or broaden—an issue that has garnered much attention Ringer:2019rfk ; Rajagopal:2016uip ; Chien:2016led ; Mehtar-Tani:2016aco ; Larkoski:2017bvj ; Chang:2017gkt ; Casalderrey-Solana:2016jvj ; Kang:2023ycg ; Kang:2023qxb ; JETSCAPE:2023hqn ; Milhano:2017nzm ; Caucal:2019uvr ; Casalderrey-Solana:2019ubu ; Wang:2019xey ; Li:2022tcr ; Budhraja:2023rgo . Measurements focused on the angular structure of inclusive jets indicate that jets become narrower in PbPb collisions at both the RHIC STAR:2021kjt and the LHC  ALargeIonColliderExperiment:2021mqf ; ATLAS:2022vii ; ALICE:2018dxf ; ALICE:2023dwg ; Ehlers:2022dfp ; ATLAS:2023hso , contrary to the theoretical expectation of intra-jet broadening Ringer:2019rfk ; Chien:2016led ; Milhano:2017nzm . In experiments, the medium modifications of the jet substructure are typically assessed by comparing two jet samples in PbPb and pp collisions that are selected within the same $p_{T}$ bins. Due to energy loss in the QGP, effectively quenched jets may be less likely to pass the $p_{T}$ selection threshold in A+A collisions, while some jets with insufficient quenching may still survive. This phenomenon is referred to as “selection bias” Baier:2001yt ; Renk:2012ve ; Wang:2021jgm ; Cunqueiro:2021wls ; Brewer:2021hmh ; Zhang:2021sua . Such biases could complicate the jet-by-jet comparison and obscure the connection between experimental measurements and intrinsic jet modifications Wang:2021jgm ; Brewer:2018dfs ; Du:2020pmp ; Brewer:2021hmh .
The V+jet, jet tagged by the vector boson ($Z^{0}/W^{\pm}$ or $\gamma$), serves as a golden channel to explore the jet quenching phenomenon in high-energy heavy-ion collisions Neufeld:2010fj ; Wang:1996yh ; Dai:2012am ; Chen:2017zte ; Wang:2013cia ; Zhang:2018urd ; Chang:2019sae ; JETSCAPE:2024rma . Since the vector boson does not interact strongly with the hot nuclear matter, it effectively gauges the initial momentum of the recoiling jet. Furthermore, the V+jet process is dominated by quark-jet production, reducing the potential influences from changes in the $q/g$ fraction during A+A collisions Wang:2020qwe . By additionally constraining the $p_{T}$ of the vector boson, the impact of selection biases on jet measurements in A+A collisions was expected to be minimized Wang:2021jgm ; Brewer:2021hmh ; STAR:2023pal ; STAR:2023ksv . Therefore, V+jet may provide unique advantages to studying medium-induced jet broadening. Recently, the CMS collaboration reports the first measurement on the two angular structure observables of $\gamma$+jet, the jet girth ($g$Larkoski:2014pca and the groomed jet radius ($R_{g}$arkoski:2014wba , in pp and PbPb collisions at $\sqrt{s}=$ 5.02 TeV CMS:2024zjn . The results show that the medium modification pattern of the angular structure significantly depends on the selection cut, $x_{j\gamma}=p_{T}^{\rm jet}/p_{T}^{\gamma}$, where $p_{T}^{\rm jet}$ and $p_{T}^{\rm\gamma}$ denote the transverse momentum of the jet and photon respectively. Notably, when setting $x_{j\gamma}>$ 0.4, there are hints of a broadening in jet angular structure at larger girth in PbPb collisions. This contrasts with the previously measured narrower girth distribution of inclusive jets reported by the ALICE collaboration ALICE:2018dxf ; ALargeIonColliderExperiment:2021mqf . The influences of selection bias in these two types of jet measurements warrant careful consideration. Timely theoretical explanations and quantitative investigations are essential for addressing this issue.
This paper presents a theoretical study of the angular structure of $\gamma$+jets in high-energy nuclear collisions at the LHC. By employing a transport approach, we carry out the medium modification of $\gamma$+jet girth in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=$ 5.02 TeV, which shows good agreement with the recently reported CMS data. Using the Jet-by-Jet matching method, we explore the relationship between selection bias and kinematic requirements in realistic event selection. Through quantitative analysis, we demonstrate that $\gamma$+jets can significantly reduce selection bias and effectively select sufficiently quenched jets in PbPb collisions compared to inclusive jets. We also discuss the influence of medium-induced gluon radiation and the medium response on the broadening of the jet angular substructure. For completeness, we will study the nuclear modification patterns of $R_{g}$ and $\Delta R_{\rm axis}$ distributions of both inclusive jets and $\gamma$+jets in PbPb collisions compared to the pp baseline.
The remainder of this paper is organized as follows. In Sec. II, we will introduce the theoretical frameworks to study the angular structure of $\gamma$+jets in pp and PbPb collisions. In Sec. III, we will discuss the main results of this paper. At last, we will summarize this study in Sec. IV.
II Theoretical framework
To investigate the angular structure of inclusive jets and $\gamma$+jets, we employ the PYTHIA8 Sjostrand:2014zea with the Monash Tune Skands:2014pea to generate the pp events as a baseline for the calculations of nucleus-nucleus collisions. Furthermore, we utilize a transport approach, which considers both the radiative and collisional partonic energy loss, to simulate the massive and massless jet evolution in the QGP. This hybrid transport approach has been used in the studies of light- and heavy-flavor dijet Dai:2018mhw ; Li:2024uzk ; Wang:2024yag ; Li:2024pfi , $Z^{0}/\gamma-$jet production in heavy-ion collisions Wang:2020qwe ; Wang:2023udp . Since the medium-induced gluon radiation plays a critical role in the jet energy loss Gyulassy:1993hr ; Wang:2001ifa , we use the radiation spectrum within the higher-twist formalism  Guo:2000nz ; Zhang:2003yn ; Zhang:2003wk ; Majumder:2009ge to simulate the in-medium jet shower in the hot/dense QCD matter,
$\frac{dN}{dxdk^{2}_{\perp}dt}=\frac{2\alpha_{s}C_{s}P(x)\hat{q}}{\pi k^{4}_{\perp}}\sin^{2}(\frac{t-t_{i}}{2\tau_{f}})(\frac{k^{2}_{\perp}}{k^{2}_{\perp}+x^{2}M^{2}})^{4},$
where $x$ and $k_{\perp}$ denote the energy fraction and transverse momentum carried by the radiated gluon. $\alpha_{s}$ is the strong coupling constant, $C_{s}$ the quadratic Casimir in color representation, $P(x)$ is the QCD splitting kernel Wang:2009qb for $q\rightarrow q+g$ and $g\rightarrow g+g$ respectively($g\rightarrow q+\bar{q}$ process is negligible due to its low possibility He:2011zx ),
$P_{q\rightarrow q+g}(x)=$ $\frac{(1-x)(1+(1-x)^{2})}{x},$ (2)
$P_{g\rightarrow g+g}(x)=$ $\frac{2(1-x+x^{2})^{3}}{x(1-x)}.$ (3)
Moreover, $\tau_{f}=2Ex(1-x)/(k^{2}_{\perp}+x^{2}M^{2})$ is the gluon formation time considering the Landau-Pomeranchuk-Migdal (LPM) effects Wang:1994fx ; Zakharov:1996fv . $\hat{q}$ denotes the jet transport parameter Chen:2010te ; Rapp:2018qla ; Cao:2018ews , determined with a $\chi^{2}$ fitting to the identified hadron production in PbPb collisions at LHC Ma:2018swx . To consider the fluctuation of medium-induced gluon radiation, we assume that the number of the radiated gluon during a time step ($\Delta t=$0.1 fm) obeys the Poisson distribution $f(n)=\lambda^{n}e^{-\lambda}/{n!}$, where the parameter $\lambda$ denotes the mean number of the radiation calculated by integrating Eq. (1). Once the radiation number $n$ is determined, the corresponding energy-momentum can be further sampled by Eq. (1) one by one. More details of the Monte Carlo implementation of the medium-induced gluon radiation can be found in Refs. Cao:2016gvr ; Wang:2020qwe .
In addition, it is also essential to consider the partonic energy loss from the elastic scattering. While the inelastic jet energy loss is carried out by the higher-twist formalisms, for completeness, the elastic energy loss is estimated by the pQCD calculation at the Hard Thermal Loop approximation Neufeld:2010xi ; Huang:2013vaa ,
$\frac{dE}{dL}=-\frac{\alpha_{s}C_{s}\mu_{D}^{2}}{2}{\rm ln}{\frac{\sqrt{ET}}{\mu_{D}}}.$ (4)
where $L$ represents the transport path along to the parton’s momentum and $\mu_{D}^{2}=6\pi\alpha_{s}T^{2}$ the Debye screening mass. The collisional energy loss of a parton can be calculated by integrating Eq. (4) during each time step. This treatment is an adequate approximation since the medium-induced gluon radiation is the dominant energy loss mechanism for light parton. The initial spacial production vertex of jets in nucleus-nucleus collisions is determined based on the MC-Glauber model Miller:2007ri . In the simulation of a jet traversing the expanding fireball, we utilize the CLVisc hydrodynamic model Pang:2016igs to generate the temperature and velocity of each medium cell. When the local temperature reaches $T_{c}=0.165$ GeV, jet parton fragment into hadron with the Colorless Hadronization prescription, which the JETSCAPE collaboration developed Putschke:2019yrg based on the Lund string model Andersson:1983jt ; Sjostrand:1984ic .
When studying jet substructures in high-energy nuclear collisions, it is important to consider the medium response effect. Energy transferred from high-$p_{T}$ jets can excite quasi-particles within the QGP medium KunnawalkamElayavalli:2017hxo ; Pablos:2019ngg ; Chen:2020tbl ; Casalderrey-Solana:2020rsj ; He:2018xjv ; Ke:2020clc . In this work, we utilize an approach based on the Cooper-Frye formula with perturbations Cooper:1974mv ; Casalderrey-Solana:2016jvj to take the medium response effect into account.
$E\frac{d\Delta N}{d^{3}p}$ $=$ $\frac{m_{T}}{32\pi T^{5}}{\rm cosh(\Delta y)exp}[-\frac{m_{T}}{T}{\rm cosh(\Delta y)]}$
×$\times\{p_{T}\Delta P_{T}{\rm cos(\Delta\phi)}+\frac{1}{3}m_{T}\Delta M_{T}{\rm cosh(\Delta y)}\}.$
Here, $\Delta y$ and $\Delta\phi$ represent the rapidity and azimuthal angle of the emitted thermal particles relative to the jet axis, respectively. Meanwhile, $m_{T}$ and $p_{T}$ denote their transverse mass and transverse momentum. The quantities $\Delta P_{T}$ and $\Delta M_{T}=\Delta E/\cosh y_{j}$ signify the transverse momentum and transverse mass transferred from the jet to the medium, where $\Delta E$ is the energy lost by the jets, and $T$ indicates the hadronization temperature of the emitted particles. Once the values for $\Delta P_{T}$ and $\Delta E$ of the jet during in-medium propagation are determined, one can sample the transverse momentum, rapidity, and azimuthal angle of the emitted particles individually based on Eq. (LABEL:eq:resp). For simplicity, we assume that all emitted partons from the excited medium hadronize into proton and pion ($\pi^{0},\pi^{\pm}$) with the ratio of $5\%$ for protons and $95\%$ for pions. Note that the above Eq. (LABEL:eq:resp) has not considered the radial flow of the expanding fireball, which cannot be neglected in the non-central nucleus-nucleus collisions. For this reason, we use the species- and momentum-dependent empirical expressions for the hadronization temperature $T$, as obtained by fitting the hadron spectra measured by the ALICE Casalderrey-Solana:2016jvj ; ALICE:2013mez .
$T_{\pi}({p_{\textrm{\tiny T}}})=\begin{cases}0.19~{}{\rm GeV},&{p_{\textrm{\tiny T}}}<0.7~{}{\rm GeV},\\ 0.21\,\left(\frac{{p_{\textrm{\tiny T}}}}{\rm{GeV}}\right)^{0.28}\,\rm{GeV},&{p_{\textrm{\tiny T}}}>0.7~{}{\rm GeV},\end{cases}$ (6)
$T_{p}({p_{\textrm{\tiny T}}})=\begin{cases}0.15~{}{\rm GeV},&{p_{\textrm{\tiny T}}}<0.07~{}{\rm GeV},\\ 0.33\,\left(\frac{{p_{\textrm{\tiny T}}}}{\rm{GeV}}\right)^{0.3}\rm{GeV},&0.07\sim 1.9~{}{\rm GeV},\\ 0.4~{}{\rm GeV},&{p_{\textrm{\tiny T}}}>1.9~{}{\rm GeV}.\end{cases}$ (7)
III Results and discussions
Recently, the medium modifications of $\gamma$+jet girth distribution in PbPb collisions $\sqrt{s_{NN}}=5.02$ TeV are measured by the CMS collaboration CMS:2024zjn . The jet girth belongs to a generalized class of observables known as jet angularities Larkoski:2014pca ; KunnawalkamElayavalli:2017hxo , defined as,
$g=\frac{1}{p_{T}^{\rm jet}}\sum_{i\in\rm jet}p_{T}^{i}\Delta R_{i,\rm jet}.$ (8)
where the index $i$ sums over all jet constituents, and $\Delta R_{i,\rm jet}$ represents the angular distance between each particle and the jet axis in the $\eta-\phi$ plane. The girth value quantifies the $p_{T}$ distribution of particles within a jet, weighted by their angular distance from the jet axis. This observable is expected to be sensitive to modifications in the jet angular structure that occur during heavy-ion collisions ALICE:2018dxf ; CMS:2024zjn ; Yan:2020zrz ; Li:2024pfi . The jet-medium interactions, such as the parton scattering, the medium-induced gluon radiation and the medium response, may increase the values of jet girth in heavy-ion collisions compared to those in pp.
Figure 1: (Color online) Normalized girth distributions of $\gamma$+jets in pp $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV calculated with $x_{j\gamma}>0.4$ (upper panel) and $x_{j\gamma}>0.8$ (middle panel) as compared to the recent CMS data CMS:2024zjn . The ratios (PbPb/pp) of girth distribution are also shown in the lower panel.
As shown in Fig. 1, we present the theoretical results of the $\gamma$+jet girth distribution for $x_{j\gamma}>0.4$ (upper panel) and $x_{j\gamma}>0.8$ (middle panel) in both pp and $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, compared with the CMS data. The lower panels show the ratio PbPb/pp distributions. All selected jets are constructed with the anti-$k_{T}$ algorithm and are required to have a transverse momentum of $p_{T}^{\rm jet}>40$ GeV, with pseudorapidity $|\eta_{\rm jet}|<2$ and jet radius $R=0.2$, while the photon must have $p_{T}^{\gamma}>100$ GeV and $|\eta_{\gamma}|<1.44$. In addition, the selected photon must meet the isolation requirement, which states that the sum of the transverse momentum of all particles within a distance $R=0.4$ around the photon must be less than 5 GeV. Furthermore, the jet and photon should be nearly “back-to-back”, requiring $\Delta\phi_{\gamma,\rm jet}>2\pi/3$. Our theoretical calculations provide a satisfactory description of the girth distribution of the CMS data for both pp and PbPb collisions for the cuts $x_{j\gamma}>0.4$ and $x_{j\gamma}>0.8$ CMS:2024zjn . Notably, a significant difference arises in the ratios of girth distributions between PbPb and pp collisions (PbPb/pp) for the two different $x_{j\gamma}$ cuts. For $x_{j\gamma}>0.4$, the modification of girth is modest, with an enhancement observed between $0.10.08<g<$ 0.1. Since girth quantifies the angular-weighted transverse momentum distribution of jets, an increase in PbPb/pp at a larger girth indicates that jets are broader in PbPb collisions compared to pp collisions. In contrast, for $x_{j\gamma}>0.8$, we observe an enhancement at $g>0.02$ and a clear suppression at $g>0.04$, suggesting that jets are narrower compared to the case of $x_{j\gamma}>0.4$. It prompts the question: how does the $x_{j\gamma}$ selection cut affect the patterns of jet angular structure modifications in nucleus-nucleus collisions? One possible explanation is that the lower $x_{j\gamma}$ cut allows for the inclusion of more significantly quenched jets traversing the QGP, which reduces the selection bias effect CMS:2024zjn . To explore this hypothesis, we conduct a Jet-by-Jet matching procedure to investigate the relationship between selection bias and kinematic cuts in heavy-ion collisions.
Figure 2: (Color online) Normalized $x_{j\gamma}$ distribution of $\gamma$+jets before and after quenching, denoted as unquenched and quenched, in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV when using different $x_{j\gamma}$ cuts by using the Jet-by-Jet matching procedure, $x_{j\gamma}>0.4$ (upper panel) and $x_{j\gamma}>0.8$ (lower panel).
In experimental measurements within the field of high-energy heavy-ion collisions, the conventional approach to study nuclear modification effects on jet substructure involves applying identical dynamical constraints during jet reconstruction and selection for both proton-proton and nucleus-nucleus collisions. By comparing the selected jet samples from these two collision systems, one can calculate the nuclear modification effects for the specific jet observables, such as jet girth. However, since the jets in these two samples do not correspond to the same individual jet before and after quenching, the derived nuclear modification effects represent an overall statistical average across a large ensemble of jets. Therefore, it is natural to inquire about how jet-medium interactions affect the substructure of individual jets before and after quenching in nucleus-nucleus collisions. While experimentally tracking the evolution of a single jet through different stages remains highly challenging, such processes can indeed be modeled and analyzed via a jet-by-jet matching in the Monte Carlo simulations as performed in the studies of Refs. Brewer:2021hmh ; Kang:2023ycg . The details are illustrated as follows:
Jet-by-Jet matching: The events selected with suitable experimental kinematic cuts, such as $p_{T}>$ 40 GeV and $|\eta_{jet}|<$ 2.0, in pp collisions are used as the input of jet evolution in PbPb. We can reconstruct the jets in pp and PbPb collisions by the event particles before and after the in-medium evolutions. By calculating the angular distance $\Delta R<R$ between the axis of each jet pair in pp and PbPb, the nearest pair of jets in the $\eta-\phi$ plane are regarded as the matched one before and after quenching. To consider the jets, with $p_{T}>$ 40 GeV initially, dropping down to the cut due to the jet energy loss, we use a lower cut $p_{T}>$ 10 GeV to select the possible candidates in PbPb collisions. This Jet-by-Jet (JBJ) matching method effectively eliminates selection bias, enabling precise tracking of the medium evolution of individual jets in the Monte Carlo simulations. It offers a novel perspective to study jet-medium interactions in high-energy heavy-ion collisions and provides critical insights for interpreting recent experimental measurements of jet substructure modifications.
As shown in Fig. 2, with the help of the Jet-by-Jet matching, we compare the $x_{j\gamma}$ distribution of $\gamma$+jets before and after quenching in $0-30\%$ PbPb collisions when using different $x_{j\gamma}$ cuts, $x_{j\gamma}>0.4$ (upper panel) and $x_{j\gamma}>0.8$ (lower panel). The solid line represents the initially selected (unquenched) jets above $x_{j\gamma}$ cut, while the dashed line denotes the corresponding jet sample after the in-medium evolution (quenched). Due to energy loss in the QGP, the $x_{j\gamma}$ distributions of quenched jets shift toward lower $x_{j\gamma}$ values compared to unquenched one. The shaded region represents the jets which have suffered sufficient quenching effect in nucleus-nucleus collisions and have dropped down to the $x_{j\gamma}$ cut. The contribution of the shaded region was usually discarded in the realistic experimental measurements in PbPb collisions, even though they initially have $p_{T}>40$ GeV. Therefore, the fraction of the shaded region quantitatively reflect the influence of selection bias, namely the amount of jets with sufficient quenching effect but are rejected by the selection requirement in the measurement in PbPb collisions. We find that the fraction of the shaded region is $23.9\%$ for $x_{j\gamma}>0.4$ and $58.4\%$ for $x_{j\gamma}>0.8$. This means that when applying the criterion $x_{j\gamma}>0.8$, less than half of the jets survive the event selection in PbPb collisions, thereby excluding a significant number of quenched jets. In contrast, using $x_{j\gamma}>0.4$ includes more jets that experience substantial quenching in PbPb collisions, ultimately resulting in a broader modification of the $\gamma$+jet girth compared to pp, as observed in the CMS measurement CMS:2024zjn .
Figure 3: (Color online) Medium modification of $\gamma$+jet girth in $0-30\%$ PbPb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV as compared to the CMS data CMS:2024zjn . In the upper panel, the results without considering the medium-induced gluon radiation and medium response are also presented. In the lower panel, we also compare the girth modification of inclusive jet and $\gamma$+jet with the same $p_{T}$ cut, as well as the case using the Jet-by-Jet matching (JBJ).
Furthermore, measurements of inclusive jets in PbPb collisions indicate a narrowing effect ALargeIonColliderExperiment:2021mqf ; ATLAS:2022vii ; ALICE:2018dxf ; ALICE:2023dwg ; Ehlers:2022dfp ; ATLAS:2023hso , while observations from $\gamma$+jets suggest hints of broadening CMS:2024zjn . It will be of great significance to theoretically explore the differing patterns of substructure modification between these two jet samples within the same collision system. In Fig. 3, we present the girth modification of $\gamma$+jets and inclusive jets in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV. Aside from the tagging requirements for $\gamma$+jets—specifically, $p_{T}^{\gamma}>100$ GeV, $x_{j\gamma}>0.4$, and $|\eta_{\gamma}|<1.44$—all selected jets must have $p_{T}>40$ GeV and $|\eta_{\text{jet}}|<2$. In the upper panel, we first discuss the influence of medium-induced gluon radiation and the medium response on $\gamma$+jets. Compared to the results without gluon radiation, our findings indicate that medium-induced gluon radiation play a significant role in the jet angular broadening in nucleus-nucleus collisions. Additionally, we also find that the medium response slightly enhances the modification in the region of larger girth. In the lower panel, we compare the girth modifications of $\gamma$+jets and inclusive jets. We observe that inclusive jets show suppression at $g>0.05$, which indicates a narrowing modification consistent with previous ALICE measurements in PbPb collisions at $\sqrt{s_{NN}}=2.76$ TeV ALICE:2018dxf . This behavior differs from that of $\gamma$+jets. To address this discrepancy, we calculate the girth modification of the jet sample using the Jet-by-Jet matching method, which represents the jet modification in the hot QCD matter without selection bias. Both $\gamma$+jets and inclusive jets show consistent and noticeable broadening at larger girth compared to the initial jets. In other words, jets naturally become broader due to jet-medium interactions in PbPb collisions for both inclusive jets and $\gamma$+jets. We will demonstrate that selection bias plays different roles in $\gamma$+jets and inclusive jets, ultimately leading to the distinct modification patterns observed in experimental measurements.
Figure 4: (Color online) Normalized $p_{T}$ distribution of unquenched and quenched jets for $\gamma$+jet and inclusive jet in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV by using the Jet-by-Jet matching procedure.
In Fig. 4, we also compare the $p_{T}$ spectra of the jets before and after quenching in PbPb collisions for both $\gamma$+jets and inclusive jets, obtained using the JBJ matching. By applying the same jet $p_{T}$ cut (specifically, $p_{T}>40$ GeV), we observe that the shape of the $\gamma$+jet $p_{T}$ spectrum before quenching significantly differs from that of the inclusive jets. The $\gamma$+jet spectrum increases gradually with $p_{T}$, peaking near the photon $p_{T}$. In contrast, the inclusive jet distribution is mainly concentrated in the range of $[40,80]$ GeV and decreases rapidly with increasing $p_{T}$. Due to in-medium energy loss, we find that the $p_{T}$ distributions of both $\gamma$+jets and inclusive jets in PbPb collisions shift towards a lower energy region compared to their pp counterparts. The shaded region indicates the jets that do not meet the $p_{T}>40$ GeV requirement due to energy loss in PbPb. Specifically, the fraction of the shaded region is $20.3\%$ for $\gamma$+jets and $63.4\%$ for inclusive jets. In other words, approximately $80\%$ of $\gamma$+jets survive the selection criteria in PbPb, whereas less than $40\%$ of inclusive jets do. This indicates that using jets associated with direct photons can significantly reduce selection bias. It can be understood from two perspectives. First, inclusive jets are mostly distributed near the selection cut (40 GeV), while jets associated with photons cover a broader $p_{T}$ range with a peak near the $p_{T}$ of the triggered photon (100 GeV). As a result, the unique $p_{T}$ spectrum of the $\gamma$+jets provides a much lower probability of jets falling below the cut-off after quenching compared to the inclusive jets. Second, the inclusive jets consist of $60\%$ gluon jets at $p_{T}^{\rm jet}>40$ GeV, whereas the $\gamma$+jet process is quark-jet dominated (above $70\%$ jets are quark-initiated jets at $p_{T}^{\rm jet}>40$ GeV for $p_{T}^{\rm\gamma}>100$ GeV). Consequently, the inclusive jet may lose more energy, resulting in a lower survival rate to pass the selection criteria in PbPb.
Figure 5: (Color online) Event averaged transverse momentum loss $\langle\Delta p_{T}\rangle=\langle p_{T}^{\rm init}-p_{T}^{\rm fin}\rangle_{\rm evt}$ of $\gamma$+jets (solid line) and inclusive jets (dash line) as a function of initial (left panel) and final (right panel) jet $p_{T}$ in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV by using the Jet-by-Jet matching procedure.
In addition to the higher survival probability of $\gamma$+jets compared to inclusive jets, we are also interested in understanding the differences in the quenching strength experienced by the surviving samples for these two types of jets within the QGP. Since quenching strength is not a specifically defined physical quantity, we assume that jet energy loss can effectively quantify the intensity of the interaction between the jet and the medium. Therefore, we estimate the event-averaged $p_{T}$ loss, denoted as $\langle\Delta p_{T}\rangle=\langle p_{T}^{\rm init}-p_{T}^{\rm fin}\rangle_{\text{evt}}$, for both $\gamma$+jets and inclusive jets as a function of final jet $p_{T}$ ($p_{T}^{\rm fin}$) in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV, using Jet-by-Jet matching as illustrated in Fig. 5. In the left panel, we plot the $p_{T}$ loss of $\gamma$+jets and inclusive jets as a function of their initial transverse momentum. It can be observed that $\langle\Delta p_{T}\rangle$ for both $\gamma$+jets and inclusive jets increases with $p_{T}^{\rm init}$, with the inclusive jets showing larger magnitudes. This aligns with expectations, as inclusive jets contain a larger fraction of gluon jets, while $\gamma$+jets are predominantly comprised of quark jets. In the right panel of Fig. 5, we also plot the $\langle\Delta p_{T}\rangle$ for $\gamma$+jets and inclusive jets as a function of $p_{T}^{\rm fin}$. Surprisingly, we find that $\langle\Delta p_{T}\rangle$ for $\gamma$+jets is significantly larger than that for inclusive jets in the range of $8040<p_{T}^{\rm fin}<$ 80 GeV. This implies that selected $\gamma$+jets with $p_{T}>40$ GeV in PbPb collisions have statistically experienced stronger quenching than the inclusive jets. Compared to inclusive jets, the specific initial $p_{T}$ distribution of $\gamma$+jets, which includes a substantial number of jets with much higher $p_{T}$ than the selection threshold, allows for the possibility that jets experiencing significant quenching can still survive in the selection in PbPb collisions. Therefore, we have quantitatively demonstrated that $\gamma$+jets provide unique and significant advantages in reducing selection bias and effectively collecting sufficiently quenched jets in PbPb collisions compared to inclusive jets.
Figure 6: (Color online) Medium modification of jet girth of $\gamma$+jets in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV relative to pp for different $x_{j\gamma}$ (0.2, 0.4, 0.6, 0.8) and $p_{T}^{\gamma}$ (80 GeV, 100 GeV, 120 GeV) cuts.
The discussions above highlight the connection between selection bias and the choice of the threshold $x_{j\gamma}=p_{T}^{\rm jet}/p_{T}^{\gamma}$. To offer theoretical guidance for searching more pronounced angular broadening effects of $\gamma$+jets in experimental settings, we have conducted calculations that explore the sensitivities of jet girth modifications to the $x_{j\gamma}$ and $p_{T}^{\gamma}$ cuts, as illustrated in Fig. 6. In the upper panel, we calculate the girth modification of $\gamma$+jet in PbPb collisions with varying $x_{j\gamma}$ cuts. It is evident that when $x_{j\gamma}>0.2$, there is more noticeable angular broadening because more quenched jets are accepted in the selection process in PbPb. As the $x_{j\gamma}$ cut increases from 0.2 to 0.8, the increasing influence of selection bias leads to a transition from enhancement at $g>0.06$ to suppression, namely the modification pattern of jet changes from broadening to narrowing, consistent with the trend observed in the CMS data CMS:2024zjn . Note that we have applied a lower $p_{T}$ cut of 20 GeV to reconstruct the jets, ensuring that the variation of $x_{j\gamma}$ down to 0.2 remains accessible. Additionally, the lower panel presents calculations of $\gamma$+jet girth modification in PbPb collisions with different $p_{T}^{\gamma}$ cuts. We find that applying a higher $p_{T}$ cut for the photon can also help mitigate the influence of selection bias and result in more pronounced girth modifications of $\gamma$+jets in PbPb collisions.
Figure 7: (Color online) Medium modification of the groomed radius distribution of $\gamma$+jet and inclusive jet in $0-30\%$ PbPb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV. The results are compared with the measurements of $\gamma$+jet by the CMS CMS:2024zjn . We also plot the results using the Jet-by-Jet matching (JBJ) as a comparison in the lower panel.
To provide a comprehensive analysis, we will also examine other jet substructure observables and highlight their different modification patterns for $\gamma$+jets and inclusive jets, similar to our findings regarding jet girth. In Fig. 7, we compare the medium modification of the groomed radius ($R_{g}$) distribution for $\gamma$+jets and inclusive jets in $0-30\%$ Pb–Pb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV. The PbPb/pp refers to the ratio of normalized $R_{g}$ distributions in Pb–Pb to pp collisions. Here, $R_{g}$ represents the opening angle between the two hardest subjets (denoted as $1$ and $2$) that satisfy the Soft Drop condition Larkoski:2014wba ,
$\frac{p_{T2}}{p_{T1}+p_{T2}}>z_{\rm cut}(\frac{R_{g}}{R})^{\beta},$ (9)
$R_{g}=\sqrt{(\eta_{1}-\eta_{2})^{2}+(\phi_{1}-\phi_{2})^{2}}.$ (10)
where $p_{T1}$ and $p_{T2}$ represent the transverse momentum of the leading and subleading subjets, respectively, while $R$ is the jet size. $\eta$ and $\phi$ are the pseudorapidity and azimuthal angle of the two subjets. We set $z_{\text{cut}}=0.2$ and $\beta=0$ in line with the CMS measurement CMS:2024zjn . The theoretical results for the modification of the groomed radius of $\gamma$+jets show a moderate enhancement at $R_{g}>0.15$, consistent with the CMS data, which may indicate jet angular broadening. In contrast, we observe that the medium modification pattern for inclusive jets shows a notable narrowing, diverging from the behavior of $\gamma$+jets. Additionally, by utilizing Jet-by-Jet matching as shown in the lower panel of Fig. 7, we can see an enhancement of the PbPb/pp ratio at $R_{g}>0.1$ for both inclusive jets and $\gamma$+jets. This observation is analogous to what we discussed earlier regarding jet girth in Fig. 3. Both inclusive jets and $\gamma$+jets become broader when they pass through the QGP. However, the selection bias reduces the broadening effect seen in the $R_{g}$ distribution of inclusive jets and $\gamma$+jets in nucleus-nucleus collisions. Despite this suppression, $\gamma$+jets exhibit a slight broadening effect, which is less influenced by selection bias compared to inclusive jets.
Figure 8: (Color online) Medium modification of the angle between jet axes of $\gamma$+jet and inclusive jet in $0-10\%$ PbPb collisions relative to pp at $\sqrt{s_{NN}}=5.02$ TeV. The results are compared with the measurements of the inclusive jet by the ALICE ALICE:2023dwg . We also plot the results using the Jet-by-Jet matching (JBJ) as a comparison in the lower panel.
In Fig. 8, we compare the medium modification of the angle between jet axes, denoted as $\Delta R_{\rm axis}$, for $\gamma$+jets and inclusive jets in $0-10\%$ PbPb collisions at $\sqrt{s_{NN}}=5.02$ TeV. The PbPb/pp represents the ratio of normalized $\Delta R_{\rm axis}$ distributions in PbPb to pp collisions. The $\Delta R_{\rm axis}$ quantifies the angular distance between the directions of jet axes reconstructed using both the standard and Winner-Take-All (WTA) recombination schemes Cal:2019gxa :
$\Delta R_{\rm axis}\equiv\sqrt{(y_{\rm axis}^{\rm WTA}-y_{\rm axis}^{\rm Std})^{2}+(\phi_{\rm axis}^{\rm WTA}-\phi_{\rm axis}^{\rm Std})^{2}}.$ (11)
Here, $y_{\rm axis}$ and $\phi_{\rm axis}$ represent the rapidity and azimuthal angle of the jet axes, respectively. The standard jet axis is derived by summing the momenta of all constituent particles of the anti-$k_{T}$ jet using a clustering algorithm known as the E-scheme. In contrast, the WTA axis is obtained by reclustering the anti-$k_{T}$ jets with the WTA recombination scheme Bertolini:2013iqa . This WTA axis is generally aligned with the most energetic constituent within the jets and minimizes the impact of soft particles on the direction of the jet axis. Therefore, the distance between the standard and WTA axes can be used to investigate the role that soft particles play in the evolution of jets in a medium. Recent studies have measured the medium modification of the inclusive jet $\Delta R_{\rm axis}$ in heavy-ion collisions, as reported by the ALICE collaboration ALICE:2023dwg . Experimental results indicate that inclusive jets show narrower $\Delta R_{\rm axis}$ distributions within the range $6040<p_{T}^{\rm jet}<$ 60 GeV in PbPb collisions compared to pp collisions. Default calculations of the $\Delta R_{\rm axis}$ modification for inclusive jets are consistent with the ALICE data. We also assess the nuclear modification of the $\gamma$+jet $\Delta R_{\rm axis}$ in the same collision system, applying different kinematic cuts ($x_{j\gamma}>0.4$, $p_{T}^{\gamma}>100$ GeV). Our calculations reveal a slight enhancement in the PbPb/pp ratio of the $\gamma$+jet $\Delta R_{\rm axis}$ distribution in PbPb collisions for $\Delta R_{\rm axis}>0.06$ compared to pp collisions. This suggests that angular jet broadening within the QGP can also be measured through the $\Delta R_{\rm axis}$ distribution of $\gamma$+jets in heavy-ion collisions. In the lower panel of Fig. 8, the Jet-by-Jet matching results exhibit clear broadening without the influence of selection bias both for $\gamma$+jet and inclusive jet, which indicates that selection bias suppresses the broadening effect of $\Delta R_{\rm axis}$ distribution in nucleus-nucleus collisions. Especially, the influence of selection bias reverses broadening to narrowing in the $\Delta R_{\rm axis}$ medium modification pattern for inclusive jets. Therefore, the differing nuclear modification patterns of $R_{g}$ and $\Delta R_{\rm axis}$ for inclusive jets and $\gamma$+jets could be of significant interest for future experimental verification. The findings presented in this paper will aid in interpreting recent measurements at the LHC and will be beneficial for future studies focusing on the intrinsic modification of jet substructure in heavy-ion collisions.
IV Summary
In summary, we present a theoretical study on the angular structure of $\gamma$+jets in high-energy nuclear collisions at the LHC. We utilize the PYTHIA8 to provide the initial production of $\gamma$+jets and employ a transport approach to simulate the in-medium jet energy loss in nucleus-nucleus collisions. We carry out the medium modification of $\gamma$+jet girth in $0-30\%$ PbPb collisions at $\sqrt{s_{NN}}=$ 5.02 TeV, which shows a satisfactory agreement with the recently reported CMS data.
Furthermore, we investigate the influence of selection bias when choosing different $x_{j\gamma}$ cuts. With the help of the Jet-by-Jet matching method, we explore the connection between selection bias and kinematic requirements in event selection. Importantly, we quantitatively demonstrate that $\gamma$+jets will provide significant advantages to reduce the selection bias and can effectively collect jets sufficiently quenched in PbPb collisions compared to the inclusive jets. We also discuss the contributions of medium-induced gluon radiation and the medium response to the broadening of jet angular substructure in PbPb collisions. Additionally, to make the discussion more general, we also study the modification patterns of $R_{g}$ and $\Delta R_{\rm axis}$ for inclusive jets and $\gamma$+jets in PbPb collisions, which show slight broadening for $\gamma$+jets but evident narrowing for inclusive jets.
This theoretical study provides new insights into the extensive measurements focusing on intra-jet broadening in heavy-ion collisions ALargeIonColliderExperiment:2021mqf ; ATLAS:2022vii ; ALICE:2018dxf ; ALICE:2023dwg ; Ehlers:2022dfp ; ATLAS:2023hso , as well as the acoplanarity broadening recently observed for lower $p_{T}$ jets STAR:2023pal ; STAR:2023ksv ; ALICE:2023qve ; ALICE:2023jye . We anticipate further measurements of $\gamma$+jet substructure in heavy-ion collisions, which may provide critical constraints on current theoretical frameworks regarding jet-medium interactions.
Acknowledgements.
This research is supported by the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008, and the National Natural Science Foundation of China with Project Nos. 11935007 and 12035007. S. W. is supported by the Open Foundation of Key Laboratory of Quark and Lepton Physics (MOE) No. QLPL2023P01 and the Talent Scientific Star-up Foundation of China Three Gorges University (CTGU) No. 2024RCKJ013.